# Retrieval-Augmented Generation for Natural Language Processing: A Survey

### v2

@@ Abstract.
Large language models (LLMs) have demonstrated great success in various fields, benefiting from their huge amount of parameters that store knowledge. However, LLMs still suffer from several key issues, such as hallucination problems, knowledge update issues, and lacking domain-specific expertise. The appearance of retrieval-augmented generation (RAG), which leverages an external knowledge database to augment LLMs, makes up those drawbacks of LLMs. This paper reviews all significant techniques of RAG, especially in the retriever and the retrieval fusions. Besides, tutorial codes are provided for implementing the representative techniques in RAG. This paper further discusses the RAG training, including RAG with/without datastore update. Then, we introduce the application of RAG in representative natural language processing tasks and industrial scenarios. Finally, this paper discusses the future directions and challenges of RAG for promoting its development.
@@ 1. Introduction
Large language models (LLMs) (Touvron et al., 2023a; Mesnard et al., 2024; Jiang et al., 2023a; OpenAI, 2023; Zeng et al., 2023) have achieved significant advancements in recent years and have become the cornerstone of various applications in the field of natural language processing (NLP). These LLMs are typically pre-trained on a large amount of natural language corpus and then fine-tuned on the specific downstream tasks’ datasets. Recent works (Petroni et al., 2019; AlKhamissi et al., 2022; Meng et al., 2022a; He et al., 2024) demonstrate the success of LLMs can be explained by the fact that language models act as knowledge bases, which refers to implicitly storing the knowledge learned from training datasets in the parameters as internal memory and generating responses by retrieving answers from memory. To store more knowledge for better generation performance, existing works generally enlarge the memory capacity by increasing the volume of parameters (Abnar et al., 2022; Brown et al., 2020; Kaplan et al., 2020; Hoffmann et al., 2022).
Although existing LLMs have shown great power, there are still several challenges hindering the development of LLMs. One of the most prominent challenges is the hallucination problem (Ji et al., 2023a; Dale et al., 2023; Ji et al., 2023b), which refers to the tendency of LLMs to generate responses that are coherent and fluent but factually incorrect. Another big challenge is the knowledge update issue. To update the knowledge stored in the LLMs’ internal memory (Meng et al., 2022a; Wang et al., 2023f; Zhang et al., 2024c), it is necessary to retrain/fine-tune LLMs with new data, which is a costly process. Another challenge for general LLMs is lacking of domain-specific expertise (Zhang et al., 2023a; Singhal et al., 2023a, b; Colombo et al., 2024). Training a domain-specific LLM, however, demands considerable manpower for dataset collection.
@@
    This paper systematically introduces each component of RAG, including details about the retriever from building to querying, and techniques of the retrieval fusions with tutorial codes.
@@
    This paper further discusses the applications of RAG on downstream NLP tasks and practical NLP scenarios.
@@
    This paper finally identifies promising future directions for exploring and main challenges for addressing.
@@
Figure 1. The overview of retrieval-augmented generation for natural language processing.
@@ 2. Overview of Retrieval-Augmented Generation
This section gives an overview of RAG for NLP. As shown in Figure 1, RAG typically consists of three modules, the retriever, the generator, and retrieval fusions.
Retriever module usually comprises three components: an encoder for encoding inputs into embeddings, an efficient indexing that supports approximate nearest neighbor search, and the datastore for storing external knowledge in the form of key-value pairs. The main challenge in the retriever module is finding the optimal trade-off between retrieval efficiency and retrieval quality. The retrieval efficiency refers to how fast the relevant information can be obtained, which involves accelerating encoding, efficient indexing, batch querying in the datastore, etc. The retrieval quality refers to how relevant the information can be retrieved, which involves chunk representation learning, advanced approximate nearest neighbor search algorithms, etc.
@@
Generator module can be classified into two branches of generators: default generators and retrieval-augmented (RA) generators. The default generators include most pre-trained/fine-tuned large language models, such as GPT-series models (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI, 2023), Mistral models (Jiang et al., 2023a), and Gemini-series models (Anil et al., 2023; Mesnard et al., 2024; Reid et al., 2024). The RA generators refer to the pre-trained/fine-tuned generators that consist of modules for fusing retrievals, such RETRO (Borgeaud et al., 2022) and Enc-Dec (Li et al., 2022a). Those generators generate responses or make predictions.
The workflow of RAG involves three steps: (1) retrieving the relevant information from external databases based on given inputs; (2) fusing the retrieved information with inputs or intermediate states based on the fusion techniques; (3) making predictions by generators based on the input and corresponding retrievals.
@@ 3.1.1. Chunking Corpus
To solve the above challenge, there are three key points that need to be considered when determining the chunking size:
@@
    Task’s property. Different tasks may benefit from different kinds of retrieval chunks. For example, question-answer tasks may prefer short phrases, while summarization tasks may prefer long documents.
@@
    Encoder’s property. Different encoder models have varying encoding capabilities on texts with different lengths. For example, models in the sentence-transformer (Reimers and Gurevych, 2019) behave better on a single sentence, while the text-embedding-ada-002 (OpenAI, 2022) is good at longer texts.
@@
    Query’s property. The length of the user’s queries should be aligned with the chunking size, which implicitly aligns the amount of contextual information in chunks with that in queries, thus improving the relevance between queries and retrievals. For example, a retrieval database built on short phrases may be useless for queries with long documents.
@@ 3.1.2. Encoding Chunks
Sparse encoding is an efficient way to encode textual chunks. However, such encodings may not capture deeper semantic meanings.
@@
    Siamese Encoders. This is a type of neural network designed to learn the similarity between inputs, which is usually trained with contrastive learning. Existing state-of-the-art siamese encoders are DPR (Karpukhin et al., 2020), SimCSE (Gao et al., 2021).
@@
    LLM-based Encoders. This type of encoder benefits from the powerful representation capability of LLMs. LLMs, which contain billions of parameters and are pre-trained on vast amounts of data covering a wide range of topics, have advanced semantic language understanding capabilities. Typical LLM-based encoders are text-embedding-ada-002 (OpenAI, 2022), bge-embedding (Xiao et al., 2023), mxbai-embedding (Sean Lee, 2024).
Compared to sparse encoding, dense encoding leverages deep neural networks, especially transformers (Vaswani et al., 2017), to capture broader linguistic and semantic information. Currently, such encodings are widely used in most semantic representation scenarios.
@@
3.1.3. Building the Index
Choice of Similarity Metrics. The similarity metrics are the basic components in the retriever, which measures the degree of relevance between query embeddings and chunk embeddings. The similarity metrics would affect the search quality. Typical similarity metrics include cosine similarity, Euclidean similarity, Manhattan distance, and Jaccard similarity.
@@ 3.1.3. Building the Index
Figure 3. The categories of fusion methods in RAG.
  1. (1)
  2. (2)
  3. (3)
3.1.4. Building the Datastore with Key-Value Pairs
The datastore used in the vector database is a specialized database that stores and manages data as a collection of key-value pairs, where keys are the unique identifier of high-dimensional embeddings and values are the domain-specific knowledge. Since the amount of the data stored in the datastore may be quite large, the storage engine, such as LMDB (LMDB, 2014) or RocksDB (Facebook, 2013), should be capable of efficient retrieval and data persistence. The key point in the datastore for ANN search is what should be used to store as values. For example, for question-answer tasks, when adding retrievals to prompts, the naive but effective way is to store the question embedding as the key and question-answer pairs as the value. This can help the generation process as retrievals are used as demonstrations for models. Recent works have proposed various state-of-the-art vector databases, including the indexing and the datastore, such as Milvus (Wang et al., 2021b; Guo et al., 2022), FAISS (Douze et al., 2024; Johnson et al., 2021), LlamaIndex (Liu, 2022), etc.
3.1.5. Code Demonstrations
3.2. Querying the Retriever
3.2.1. Encoding Queries and ANN Search
To align with the pre-built embedding space, the retriever uses the same encoder to encode queries during the querying stage. The ANN search refers to leveraging the pre-built indexing and datastore to find similar data via approximate nearest neighbor searching algorithms and then retrieve the corresponding values.
Retrieving values from datastore refers to fetching the corresponding values based on the key identifiers of nearest neighbors.
@@ 4.2. Logits-based Fusion
6:        $h^{R}_{i}=\mathcal{M}^{C}_{i}(h^{A}_{i},e_{v_{1}},\ldots,e_{v_{k}})$; /* Use a cross-attention module to incorporate external knowledge */
@@
15:        $h^{R}_{i}=h^{A}_{i}+\frac{1}{k}\sum_{j}w_{j}e_{v_{j}}$ /* Use a weighted sum mechanism to fuse the retrieved knowledge */
@@ 4.3. Latent Fusion
5. Generators
This section introduces representative generators and retrieval-augmented generators, which are generally pre-trained on large datasets. Existing generators are mostly large language models that adopt or modify the transformer-based architecture (Vaswani et al., 2017). For example, Llama-series models (Touvron et al., 2023a; Touvron et al., 2023b), GPT-series models (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI, 2023), and Gemini-series models (Anil et al., 2023; Reid et al., 2024; Mesnard et al., 2024) remove all encoder modules, retaining only the decoder module, which includes an attention module and a feed-forward network module. Other advanced techniques, such as root mean square layer normalization (Zhang and Sennrich, 2019), rotary position embedding (Su et al., 2024), and group query attention mechanisms (Ainslie et al., 2023), have been incorporated into the design of existing large language models to enhance their performance.
Retrieval-augmented generators typically incorporate new modules into the architecture of existing large language models. They are also pre-trained on a large dataset and an external knowledge database constructed from a vast natural language corpus. These generators mostly leverage latent fusions to incorporate the knowledge into the hidden states of large language models (Borgeaud et al., 2022; Wu et al., 2022; Li et al., 2022a), which has been discussed in Section 4.3.
6. RAG Training
@@ 6.1. RAG without Datastore Update
Considering the case of no datastore update, training the retriever generally refers to training the retriever encoder and rebuilding the indexing. Since sparse encodings typically rely on statistical methods without parameters, training the encoder pertains only to dense encoding methods. Different training methods may have different goals, such as improving the semantic representations, accelerating the encoding process, or learning the domain-specific representations. The first two goals are often achieved by replacing the original encoder with a more powerful or tiny encoder, such as DistilBERT (Sanh et al., 2019) or TinyBERT (Jiao et al., 2020). The last requires training the original encoder on the domain-specific corpus using contrastive learning. After training the retriever encoder, the embeddings that serve as keys in the vector database will also change. Thus, all indexes should be rebuilt with new embeddings. Besides, if the encoder remains unchanged, the indexing can be updated using new ANN searching algorithms or re-tuning the hyperparameters. After the retriever is trained, it can be directly incorporated into the RAG without updating the generator.
@@
Apart from independently training the retriever and the generator, jointly training the retriever and the generator can be another good choice for better performance on downstream tasks. The key to this case is to ensure the differentiability from the input to the output during the forward process. Typically, complex indexes, such as FAISS (Douze et al., 2024), are not a suitable choice during the fine-tuning stage. Existing works generally leverage the complex indexes to pre-select a small subset of nearest neighbors as candidates, then choose the final top-$k$ nearest neighbors by performing the matrix-multiplication operations. Joint training is an end-to-end optimization that can lead to better coordination between the retriever and the generator and improve the contextual understanding of the generator.

### v3

@@ Abstract.
Large language models (LLMs) have demonstrated great success in various fields, benefiting from their huge amount of parameters that store knowledge. However, LLMs still suffer from several key issues, such as hallucination problems, knowledge update issues, and lacking domain-specific expertise. The appearance of retrieval-augmented generation (RAG), which leverages an external knowledge database to augment LLMs, makes up those drawbacks of LLMs. This paper reviews all significant techniques of RAG, especially in the retriever and the retrieval fusions. Besides, tutorial codes are provided for implementing the representative techniques in RAG. This paper further discusses the RAG update, including RAG with/without knowledge update. Then, we introduce RAG evaluation and benchmarking, as well as the application of RAG in representative NLP tasks and industrial scenarios. Finally, this paper discusses RAG’s future directions and challenges for promoting this field’s development.
Reference Format:
*Corresponding author.
@@ 1. Introduction
Large language models (LLMs) (Touvron et al., 2023a; Mesnard et al., 2024; Jiang et al., 2023a; OpenAI, 2023; Zeng et al., 2023) have achieved significant advancements in recent years and have become the cornerstone of various applications in the field of natural language processing (NLP). These LLMs are typically pre-trained on a large amount of natural language corpus and then fine-tuned on the specific downstream tasks’ datasets. Recent works (Petroni et al., 2019; AlKhamissi et al., 2022; Meng et al., 2022a; He et al., 2024) demonstrate the success of LLMs can be explained by the fact that LLMs act as knowledge bases, which refers to implicitly storing the learned knowledge in the parameters as internal memory and generating responses by retrieving answers from memory. To store more knowledge for better generation performance, existing works generally enlarge the memory capacity by increasing the volume of parameters (Abnar et al., 2022; Brown et al., 2020; Kaplan et al., 2020; Hoffmann et al., 2022).
Although existing LLMs have shown great power, several challenges still hinder the development of LLMs. One of the most prominent challenges is the hallucination problem (Ji et al., 2023a; Dale et al., 2023; Ji et al., 2023b), which refers to the tendency of LLMs to generate responses that are coherent and fluent but factually incorrect. Another big challenge is the knowledge update issue. To update the knowledge stored in the LLMs’ internal memory (Meng et al., 2022a; Wang et al., 2023g; Zhang et al., 2024c), it is necessary to retrain/fine-tune LLMs with new data, which is a costly process. Another challenge for general LLMs is lacking domain-specific expertise (Zhang et al., 2023a; Singhal et al., 2023a, b; Colombo et al., 2024). Training a domain-specific LLM demands considerable manpower for dataset collection.
@@
Figure 1. The overview of retrieval-augmented generation for natural language processing. The inputs as queries are fed into both the retriever for retrieval knowledge and the generator for outputs. There are three kinds of retrieval fusions, including query-based fusion, logits-based fusion, and latent fusion.
[…]
@@
    This paper systematically introduces each component of RAG, including details about the retriever from building to querying and techniques of the retrieval fusions with tutorial codes.
@@
    This paper further discusses RAG evaluation and benchmarking, as well as the applications of RAG on downstream NLP tasks and practical NLP scenarios.
@@
    This paper finally identifies promising future directions for exploring and the main challenges for addressing.
@@ 2. Overview of Retrieval-Augmented Generation
This section gives an overview of Retrieval-Augmented Generation (RAG) for NLP. An RAG system utilizes the external knowledge base $\boldsymbol{D}$ to enhance the generation system. Taking external documents as an example, $\boldsymbol{D}$ consists of external documents, each of which contains a set of chunks $c_{i}\in\boldsymbol{C_{i}}$. These chunks are transformed into vector embedding using an embedding model. When inputting a query $q$, which is embedded as a vector, then the retriever in the RAG system retrieves top-$k$ chunks $R_{q}=\{r_{1},r_{2},\dots,r_{k}\}$ most relevant to the query $q$. The RAG system can use different fusion methods to fuse the retrieved chunks, which is discussed in Section 4. The overall process is formulated as follows:
(1) $\texttt{Retriever}(q,D)\rightarrow R_{q}$
(2) Generator(Retrieval $\texttt{Generator}(\texttt{Retrieval Fusion}(q,R_{q}))\rightarrow answer$
As shown in Figure 1, RAG typically consists of three modules, the retriever, the generator, and retrieval fusions.
Retriever module usually comprises three components: an encoder for encoding inputs into embeddings, an efficient indexing that supports approximate nearest neighbor search, and a datastore for storing external knowledge in the form of key-value pairs. The main challenge in the retriever module is finding the optimal trade-off between retrieval efficiency and retrieval quality. The retrieval efficiency refers to how fast the relevant information can be obtained, which involves accelerating encoding, efficient indexing, batch querying in the datastore, etc. The retrieval quality refers to how relevant the information can be retrieved, which involves chunk representation learning, advanced approximate nearest neighbor search algorithms, etc.
@@
Generator modules can be classified into two branches of generators: default generators and retrieval-augmented (RA) generators. The default generators include most pre-trained/fine-tuned large language models, such as GPT-series models (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI, 2023), Mistral models (Jiang et al., 2023a), and Gemini-series models (Anil et al., 2023; Mesnard et al., 2024; Reid et al., 2024). The RA generators refer to the pre-trained/fine-tuned generators that consist of modules for fusing retrievals, such as RETRO (Borgeaud et al., 2022; Wang et al., 2023e) and Enc-Dec (Li et al., 2022a). Those generators generate responses or make predictions.
The workflow of RAG involves three steps: 1) retrieving the relevant information from external databases based on given inputs; 2) fusing the retrieved information with inputs or intermediate states based on the fusion techniques; 3) making predictions by generators based on the input and corresponding retrievals.
@@ 3.1.1. Chunking Corpus
To solve the above challenge, three key points need to be considered when determining the chunking size:
@@
    Task’s preference. Different tasks may benefit from different kinds of retrieval chunks. For example, question-answer tasks may prefer short phrases, while summarization tasks may prefer long documents.
@@
    Encoder’s preference. Different encoder models have varying encoding capabilities on texts with different lengths. For example, models in the sentence-transformer (Reimers and Gurevych, 2019) behave better on a single sentence, while the text-embedding-ada-002 (OpenAI, 2022) is good at longer texts.
@@
    Query’s preference. The length of the user’s queries should be aligned with the chunking size, which implicitly aligns the amount of contextual information in chunks with that in queries, thus improving the relevance between queries and retrievals. For example, a retrieval database built on short phrases may be useless for queries with long documents.
@@ 3.1.2. Encoding Chunks
  3. (3)
    BM25 (Robertson and Zaragoza, 2009) is a probabilistic ranking algorithm used in information retrieval to estimate the relevance of documents to a search query by balancing term frequency, inverse document frequency, and document length normalization, ensuring robust scoring even for long or short documents. BM25 focuses on lexical matches and is computationally efficient, making it a cornerstone of traditional search engines.
Sparse encoding is an efficient way to encode textual chunks. However, such encoding methods may not capture deeper semantic meanings well.
@@
    Siamese Encoders. This is a type of neural network designed to learn the similarity between inputs, which is usually trained with contrastive learning. Existing state-of-the-art siamese encoders are DPR (Karpukhin et al., 2020), SimCSE (Gao et al., 2021), Contriever (Izacard et al., 2022).
@@
    LLM-based Encoders. This type of encoder benefits from the powerful representation capability of LLMs. LLMs, which contain billions of parameters and are pre-trained on vast amounts of data covering a wide range of topics, have advanced semantic language understanding capabilities. Typical LLM-based encoders are text-embedding-ada-002 (OpenAI, 2022), bge-embedding (Xiao et al., 2023), mxbai-embedding (Sean Lee, 2024), MedCPT (Jin et al., 2023).
Compared to sparse encoding, dense encoding leverages deep neural networks, especially transformers (Vaswani et al., 2017), to capture broader linguistic and semantic information. Dense encodings are widely used in most representation scenarios. Moreover, some works also utilized hybrid methods to encode text for leveraging both lexical and semantic information (Lewis et al., 2022; Luo et al., 2023).
3.1.3. Building the Index
Choice of Similarity Metrics. The similarity metrics are the basic components in the retriever, which measures the degree of relevance between query embeddings and chunk embeddings. The similarity metrics would affect the search quality. Typical similarity metrics include cosine similarity, Euclidean similarity, and Manhattan distance.
  1. (1)
  2. (2)
  3. (3)
3.1.4. Building the Datastore with Key-Value Pairs
The datastore used in the vector database is a specialized database that stores and manages data as a collection of key-value pairs, where keys are the unique identifiers of high-dimensional embeddings and values are the domain-specific knowledge. Since the amount of the data stored in the datastore may be quite large, the storage engine, such as LMDB (LMDB, 2014) or RocksDB (Facebook, 2013), should be capable of efficient retrieval and data persistence. The key point in the datastore for ANN search is what should be used to be stored as values. For example, for question-answer tasks, when adding retrievals to prompts, the naive but effective way is to store the question embedding as the key and question-answer pairs as the value. This can help the generation process as retrievals are used as demonstrations for models. Recent works have proposed various state-of-the-art vector databases including the indexing and the datastore, such as Milvus (Wang et al., 2021b; Guo et al., 2022), FAISS (Douze et al., 2024; Johnson et al., 2021), LlamaIndex (Liu, 2022), etc.
@@ 3.1.4. Building the Datastore with Key-Value Pairs
3.1.5. Code Demonstrations
3.2. Querying the Retriever
3.2.1. Encoding Queries and ANN Search
To align with the pre-built embedding space, the retriever uses the same encoder to encode queries during the querying stage. The ANN search leverages the pre-built indexing and datastore to find similar data via ANN searching algorithms and then retrieves the corresponding values.
Retrieving values from datastore fetches the corresponding values based on the nearest key identifiers.
@@ 4.2. Logits-based Fusion
6:        $h^{R}_{i}=\mathcal{M}^{C}_{i}(h^{A}_{i},e_{v_{1}},\ldots,e_{v_{k}})$; /* Cross-attention */
@@
15:        $h^{R}_{i}=h^{A}_{i}+\frac{1}{k}\sum_{j}w_{j}e_{v_{j}}$ /* Weighted sum */
@@ 4.3. Latent Fusion
4.4. Comparison of Different Fusion Methods
Each fusion method exhibits distinct advantages and limitations. The query-based fusion method is a simple and straightforward approach that preserves the raw information from retrieved information. However, it significantly increases the input sequence length, resulting in higher computational costs. Additionally, this kind of method suffers from limited contextual integration capabilities and scalability, as it struggles to effectively combine retrieved information with the input context.
The logits-based fusion method combines the output logits (or probabilities) of the base LLM with those from a separate model (or shared LLM) that processes the retrieved information. This approach is computationally efficient, as it can process retrievals and inputs in batches, avoiding increasing the input sequence length. It decouples the retrieval and generation processes, enabling independent improvements to each component. Nevertheless, this kind of method is limited by its shallow contextual integration, as fusion occurs only at the output level. Consequently, it may underperform in complex reasoning tasks that require fine-grained interaction between the input and retrieved information.
In contrast, the latent fusion method integrates retrieved information directly into the hidden states (or intermediate representations) of the LLM. This approach is both scalable and efficient, as it avoids expanding the input sequence length. Moreover, it facilitates deep, context-aware integration of retrieved information and supports fine-grained control over how the information is fused. However, the latent fusion method necessitates significant architectural modifications and careful training to achieve optimal performance. In summary, the choice of fusion method depends on the specific application, available resources, and the desired trade-off between simplicity and performance. Each method offers unique strengths and weaknesses, making them suitable for different use cases and operational constraints.
5. Generators
This section introduces representative generators, which are generally pre-trained on large datasets. Existing generators are mostly large language models that adopt or modify the transformer-based architecture (Vaswani et al., 2017). For example, DeepSeek (Guo et al., 2025), Llama-series models (Touvron et al., 2023a, b), GPT-series models (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI, 2023), and Gemini-series models (Anil et al., 2023; Reid et al., 2024; Mesnard et al., 2024) remove all encoder modules, retaining only the decoder module, which includes an attention module and a feed-forward network module. Other advanced techniques, such as root mean square layer normalization (Zhang and Sennrich, 2019), rotary position embedding (Su et al., 2024), and group query attention mechanisms (Ainslie et al., 2023), have been incorporated into the design of existing large language models to enhance their performance.
Existing generators can be generally categorized into two groups, close-sourced LLMs (Brown et al., 2020; OpenAI, 2023) and open-sourced LLMs (Guo et al., 2025; Touvron et al., 2023a, b). The formers can only adopt query-based fusions to introduce external knowledge. While the latter can adapt to all kinds of fusions. Especially generators using the latent fusion typically would train the generators from scratch. Those generators integrate novel modules designed for retrieved information fusion, such as cross-attention-based module (Borgeaud et al., 2022; Wu et al., 2022; Li et al., 2022a). Their training paradigm combines pre-training on massive textual datasets with integrating external retrieved knowledge to establish robust information grounding. As detailed in Section 4.3, those generators enable dynamic, context-aware retrieved information incorporation while further improving the LLM’s generative capabilities.
6. RAG Training and Datastore Update
@@
[…]
@@ 6.1. RAG without Datastore Update
Considering the case of no datastore update, training the retriever generally refers to training the retriever encoder and rebuilding the indexing. Since sparse encodings rely on statistical methods without parameters, training the encoder pertains only to dense encoding methods. Different training methods may have different goals, such as improving the semantic representations, accelerating the encoding process, or learning the domain-specific representations. The first two goals are often achieved by replacing the original encoder with a more powerful or tiny encoder, such as DistilBERT (Sanh et al., 2019), or TinyBERT (Jiao et al., 2020). The last requires training the original encoder on the domain-specific corpus. REPLUG (Shi et al., 2023) updated the retriever by minimizing KL divergence between retrieval and LM scores, and asynchronously refreshing the datastore index. After training the retriever encoder, the vector database’s embeddings that serve as keys will also change. Thus, all indexes should be rebuilt with new embeddings. Besides, if the encoder remains unchanged, the indexing can be updated using new ANN searching algorithms or re-tuning the hyperparameters. After the retriever is trained, it can be directly incorporated into the RAG without updating the generator.
@@
Apart from independently training the retriever and the generator, jointly training the retriever and the generator can be another good choice for better performance on downstream tasks. The primary challenges involve computational complexity from large-scale retrieval and marginalization over latent documents, alongside training instability due to potential retrieval collapse and interdependent retriever-generator feedback loops. Additionally, sparse gradients from discrete retrieval steps and approximation errors from fixed top-$k$ document selection limit effective end-to-end optimization. Typically, complex indexes, such as FAISS (Douze et al., 2024), are not a suitable choice during the fine-tuning stage.
Existing works generally leverage the complex indexes to pre-select a small subset of nearest neighbors as candidates, then choose the final top-$k$ nearest neighbors by performing the matrix-multiplication operations. REALM (Guu et al., 2020) and RAG (Lewis et al., 2020) both unify retriever and language modeling by treating retrieved documents as latent variables and optimizing end-to-end via gradient descent. They both used Maximum Inner Product Search (MIPS). RAG (Lewis et al., 2020) only updated the encoder in the retriever, but REALM (Guu et al., 2020) also asynchronous refreshed the indexing. Atlas (Izacard et al., 2023) jointly pre-trained the retriever and generator, using different loss functions (such as attention distillation and perplexity distillation) to optimize the retriever. Besides, it also asynchronous refreshed the indexing. Experimental results in these works demonstrate that joint training is an end-to-end optimization that can lead to better coordination between the retriever and the generator and improve the contextual understanding of the generator.